# Sparse but Wrong: Incorrect L0 Leads to Incorrect Features in Sparse Autoencoders

### v2

@@ 1 Introduction
It is theorized that Large Language Models (LLMs) represent concepts as linear directions in representation space, known as the Linear Representation Hypothesis (LRH) (Elhage et al., 2022; Park et al., 2024). These concepts are nearly orthogonal linear directions, allowing the LLM to represent many more concepts than there are neurons, a phenomenon known as superposition (Elhage et al., 2022). However, superposition poses a challenge for interpretability, as neurons in the LLM are polysemantic, firing on many different concepts.
@@
Sparse autoencoders (SAEs) are meant to reverse superposition, and extract interpretable, monosemantic latent features (Cunningham et al., 2024; Bricken et al., 2023) using sparse dictionary learning (Olshausen & Field, 1997). SAEs have the advantage of being unsupervised, and can be scaled to millions of neurons in its hidden layer (hereafter called “latents” [^1] We use latents to prevent overloading the term feature, which we reserve for human-interpretable concepts the SAE may capture. This breaks from earlier usage which used feature for both (Elhage et al., 2022), but aligns with the terminology in (Lieberum et al., 2024) and makes the distinction more clear.). When training an SAE, practitioners must decide on the sparsity of SAE, measured in terms of L0, or how many latents activate on average for a given input.  [^2] TopK and BatchTopK SAEs (Gao et al., 2024; Bussmann et al., 2024) set the L0 ($K$) directly, whereas L1 and JumpReLU (Cunningham et al., 2024; Bricken et al., 2023; Rajamanoharan et al., 2024) adjust it via a coefficient in the loss. In any case, all SAE trainers must decide on the target L0. L0 is typically considered a neutral design choice: most of the literature evaluates SAEs at a range of L0 values, referring to this as a “sparsity–reconstruction tradeoff” (Gao et al., 2024; Rajamanoharan et al., 2024). While most practitioners would expect that too high an L0 will break the SAE (afterall, it is called a sparse autoencoder), the implication of “sparsity–reconstruction tradeoff” plots is that any sufficiently low L0 is equally valid.
@@
However, recent work shows the same trend: low L0 SAEs perform worse on downstream tasks Kantamneni et al. (2025); Bussmann et al. (2025). What causes this degraded performance at low L0? In this work, we explore the effect of L0 on SAEs. We begin with toy model experiments using synthetic data, and show that if the L0 is too low, the SAE can “cheat” by mixing together components of correlated features, achieving better reconstruction compared to an SAE with correct, disentangled features. We consider this to be a manifestation of feature hedging (Chanin et al., 2025), where the SAE abuses feature correlations to compensate for insufficient resources to model the underlying features monsemantically. This mixing of correlated features into SAE latents affects both positively and negatively correlated features, meaning that in low L0 SAEs, nearly all latents are both less interpretable and more noisy than an SAE with a correctly set L0.
@@
Finally, we develop a proxy metric based on projections between the SAE decoder and training activations that can detect if L0 is too low. We validate these findings on Gemma-2-2b (Team et al., 2024), demonstrating that decoder patterns similar to what we observe in our toy model experiments also manifests in LLM SAEs. We further validate that the optimal L0 we find with our method in Gemma-2-2b matches peak performance on sparse probing tasks (Kantamneni et al., 2025).
@@ 2 Background
An SAE decomposes an input activation $\mathbf{x}\in\mathbb{R}^{d}$ into a hidden state, $\mathbf{a}$, consisting of $h$ hidden neurons, called “latents”. An SAE is composed of an encoder $\mathbf{W}_{\text{\text{enc}}}\in\mathbb{R}^{h\times d}$, a decoder $\mathbf{W}_{\text{\text{dec}}}\in\mathbb{R}^{d\times h}$, a decoder bias $\mathbf{b}_{\text{\text{dec}}}\in\mathbb{R}^{d}$, and encoder bias $\mathbf{b}_{\text{\text{enc}}}\in\mathbb{R}^{h}$, and a nonlinearity $\sigma$, typically ReLU or a variant like JumpReLU (Rajamanoharan et al., 2024), TopK (Gao et al., 2024) or BatchTopK (Bussmann et al., 2024). The decoder is sometimes called the dictionary, in reference to sparse dictionary learning. We use both terms interchangeably.
@@
In this work we focus on BatchTopK and JumpReLU SAEs as these are both considered SOTA architectures. The JumpReLU activation is a modified ReLU with a threshold parameter $\tau>0$, so $\text{JumpReLU}_{\tau}(x)=x\cdot\mathbf{1}_{x>\tau}$. The BatchTopK activation function selects the top $b\times k$ activations across a batch of size $b$, allowing variance in the $k$ selected per sample in the batch. After training, a BatchTopK SAE is converted to a JumpReLU SAE with a global $\tau$. We follow the JumpReLU training procedure outlined by Anthropic (Conerly et al., 2025).
@@ 3 Toy model experiments
The Linear Representation Hypothesis (LRH) (Elhage et al., 2022; Park et al., 2024) states that LLMs represent concepts (alternatively referred to as “features”) as (nearly) orthogonal linear directions in representation space. Thus, the hidden activations in an LLM are simply the sum all the firing feature vectors (a feature direction with a positive, non-zero magnitude) that are being represented. While an LLM can represent a potentially large number of concepts this way, in any given activation, only a small number of concepts are actively represented.
@@
In these toy model experiments, we mainly focus on BatchTopK SAEs (Bussmann et al., 2024) as this enables direct control of L0. Additionally, we validate our results with JumpReLU SAEs. We train SAEs on 15M synthetic samples with batch size 500 using SAELens (Bloom et al., 2024).
@@ 3.4 The sparsity–reconstruction tradeoff
It is common practice to evaluate SAE architectures using a sparsity–reconstruction tradeoff plot (Cunningham et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024), where the assumption is that having better reconstruction at a given sparsity is inherently better, and indicates that the SAE is correct. Afterall, we train SAEs to reconstruct inputs, so surely an SAE that has better reconstruction must therefore be a better SAE than one that has lower reconstruction?
Sadly, this is not the case. As we discussed in Section 3.3, when the L0 of the SAE is lower than optimal, the SAE can find ways to “cheat” by engaging in feature hedging (Chanin et al., 2025), and get a better MSE score by mixing components of correlated features together. This results in an SAE where the latents are not monosemantic, and do not track ground-truth features.
@@
Figure 5: SAE decoder cosine similarity with true features for the learned SAEs with L0=1 (left) and L0=2 (middle), compared with the ground-truth SAE (right). The learned SAEs score much better than the ground truth SAE on variance explained, despite their corrupted, polysemantic latents.
@@ 3.5 Detecting the true L0 using the SAE decoder
Figure 6: Idealized histogram of decoder projections on input activations demonstrating the intuition behind our $n$th decoder projection metric, $s_{n}^{\text{dec}}$. For an arbitrary input, most latents should be non-active and thus have low projection. When SAE latents are monosemantic, we expect non-active latents to have a near-zero projection on arbitrary input activations. However, if SAE latents mix positive and negative components of many underlying features, then those latents will have larger projections on arbitrary inputs that contains those features. By picking an NN less than $h/2$, a smaller $s_{n}^{\text{dec}}$ means latents have smaller projection on arbitrary inputs and thus are more monosemantic.
$s_{n}^{\text{\text{dec}}}=\mathbf{z}_{\downarrow}[n\cdot b]$ (5)
where $n\in\{1,2,\ldots,h\}$ indexes the latent by its ranking. The multiplication by $b$ accounts for the batch dimension, effectively selecting the $n$th highest projection value when considering all samples in the batch. For this to work $n$ should be sufficiently larger than a reasonable guess at the correct L0, as in a perfect SAE, the decoder for these latents should be uncorrelated with input activations. Empirically, picking $n$ up to about $\nicefrac{{h}}{{2}}$ seems to work well.
Alternative metric: decoder pairwise cosine similarity
We find that $s_{n}^{\text{\text{dec}}}$ is not the only metric that works to locate the ideal L0. Taking the mean cosine similarity between every latent in the decoder also works well. We focus on $s_{n}^{\text{\text{dec}}}$ as it allows for more interesting analysis, but we do not feel either metric is inherently better than the other, and recommend practitioners experiment with both. We explore this alternative metric further in Appendix A.5.
@@ 3.6 JumpReLU SAE experiments
So far, we have only investigated BatchTopK SAES due to their ease of setting L0. We now validate that these same conclusions apply to JumpReLU SAEs. We train JumpReLU saes with a range of $\lambda_{s}$ to control the sparsity of the SAEs. We show plots of $\lambda_{s}$ vs L0 $n$th decoder projection vs L0 for these SAEs in Figure 8. We see that $n$th decoder projection vs L0 broadly follows the same pattern as we saw for BatchTopK SAEs, and is minimized at and slightly above the correct L0.
Figure 8: (left) L0 coefficient $\lambda_{s}$ vs L0 for JumpReLU SAEs. (right) $n$th decoder projection vs L0 for JumpReLU SAEs. The true L0, 11, is marked by a dotted line on the plot.
Interestingly, we see that the L0 does not change linearly with $\lambda_{s}$, but instead “sticks” near the correct L0. This is a testament to Anthropic’s JumpReLU SAE training method (Conerly et al., 2025), as a wide range of sparsity coefficients $\lambda_{s}$ cause the SAE to naturally find the correct L0.
@@ 4 LLM experiments
We train a series of BatchTopK SAEs (Bussmann et al., 2024) with $h=32768$ on Gemma-2-2b (Team et al., 2024) and Llama-3.2-1b (Dubey et al., 2024) varying L0 and calculate $s_{n}^{\text{\text{dec}}}$. Each SAE is trained on 500M tokens from the Pile (Gao et al., 2020) using SAELens (Bloom et al., 2024). We also calculate k-sparse probing performance for these SAEs using the benchmark from Kantamneni et al. (2025), consisting of over 100 sparse probing tasks. Results are shown in Figure 9 for $n=16000$. We find that using $n$ near $h/2$ seems to achieve best results in our LLM experiments.
Figure 9: $n$ = 16k decoder projection vs SAE L0 and K-sparse probing F1 vs L0 with 3 seeds per L0. (left) Gemma-2-2b layer 5 BatchTopK results. (right) Llama-3.2-1b layer 7 BatchTopK SAEs. In both cases, peak sparse probing performance occurs in the elbow just before $s_{n}^{\text{\text{dec}}}$ jumps due to low L0, although the shapes of the $s_{n}^{\text{\text{dec}}}$ plots vary at high L0.
The Llama SAE $s_{n}^{\text{dec}}$ plot looks very similar to the toy model, with a clear minimum point. The Gemma-2-2b layer 5 SAEs also show a sharp increase in $s_{n}^{\text{dec}}$ at low L0 as we saw in toy models, but has a long shallow region with the global minimum actually appearing in that shallow region. In both cases, the “elbow” in the $s_{n}^{\text{dec}}$ plots just before the jump due to low L0 is around L0 200, and this also corresponds to peak sparse probing performance. More plots and analysis of $s_{n}^{\text{dec}}$ curves are shown in Appendix A.9.
Figure 10: Gemma-2-2b layer 12, with (left) $n$th decoder projection and K-sparse probing F1 for BatchTopK and JumpReLU SAEs, and (right) normalized decoder projection histograms for BatchTopK SAEs. The histograms are truncated to -20 and 40 to highlight projections near the origin.
@@ 4.1 JumpReLU vs BatchTopK SAEs
We next explore how JumpReLU and BatchTopK SAEs compare with $n$th decoder projection plots. We train a suite of SAEs on 1B tokens on Gemma-2-2b layer 12. We plot $s_{n}^{\text{dec}}$ for a range N values as well as k-sparse probing results for JumpReLU and BatchTopK SAEs in Figure 10 (left).
JumpReLU and BatchTopK SAEs behave similarly at low L0, with the high $s_{n}^{\text{dec}}$ at low L0 corresponding to poor sparse-probing performance. However, we see notable differences at high L0. The BatchTopK SAEs have a global $s_{n}^{\text{dec}}$ minimum around 200-250 no matter the N value, but JumpReLU SAEs only have a clear $s_{n}^{\text{dec}}$ minimum at 200-250 when $n=16000$ (about $h/2$). As we saw in Figure 9 as well, using the “elbow” of the plots just before $s_{n}^{\text{dec}}$ jumps due to low L0 seems to correspond to peak k-sparse probing performance.
For JumpReLU SAEs, we see that $s_{n}^{\text{dec}}$ does not rise much at high L0, and indeed, JumpReLU SAEs also perform much better than BatchTopK SAEs at sparse probing when L0 is high. We suspect this is due to JumpReLU SAEs being able to “stick” near the correct threshold per latent like we saw in our toy models section. We investigate the differences in learned SAEs between JumpReLU and BatchTopK further in Appendix A.10.

### v3

@@ 1 Introduction
It is theorized that Large Language Models (LLMs) represent concepts as linear directions in representation space, known as the Linear Representation Hypothesis (LRH) (elhage2022toy; park2024lrh). These concepts are nearly orthogonal linear directions, allowing the LLM to represent many more concepts than there are neurons, a phenomenon known as superposition (elhage2022toy). However, superposition poses a challenge for interpretability, as neurons in the LLM are polysemantic, firing on many different concepts.
@@
Sparse autoencoders (SAEs) are meant to reverse superposition, and extract interpretable, monosemantic latent features (huben2024sparse; bricken2023towards) using sparse dictionary learning (olshausen1997sparse). SAEs have the advantage of being unsupervised, and can be scaled to millions of neurons in its hidden layer (hereafter called “latents” [^1] We use latents to prevent overloading the term feature, which we reserve for human-interpretable concepts the SAE may capture. This breaks from earlier usage which used feature for both (elhage2022toy), but aligns with the terminology in (lieberumGemmaScopeOpen2024) and makes the distinction more clear.). When training an SAE, practitioners must decide on the sparsity of SAE, measured in terms of L0, or how many latents activate on average for a given input.  [^2] TopK and BatchTopK SAEs (gao2024scaling; bussmann2024batchtopk) set the L0 ($K$) directly, whereas L1 and JumpReLU (huben2024sparse; bricken2023towards; rajamanoharan2024jumping) adjust it via a coefficient in the loss. In any case, all SAE trainers must decide on the target L0. L0 is typically considered a neutral design choice: most of the literature evaluates SAEs at a range of L0 values, referring to this as a “sparsity–reconstruction tradeoff” (gao2024scaling; rajamanoharan2024jumping). While most practitioners would expect that too high an L0 will break the SAE (afterall, it is called a sparse autoencoder), the implication of “sparsity–reconstruction tradeoff” plots is that any sufficiently low L0 is equally valid.
@@
However, recent work shows the same trend: low L0 SAEs perform worse on downstream tasks kantamneni2025sparse; bussmann2025learning. What causes this degraded performance at low L0? In this work, we explore the effect of L0 on SAEs. We begin with toy model experiments using synthetic data, and show that if the L0 is too low, the SAE can “cheat” by mixing together components of correlated features, achieving better reconstruction compared to an SAE with correct, disentangled features. We consider this to be a manifestation of feature hedging (chanin2025feature), where the SAE abuses feature correlations to compensate for insufficient resources to model the underlying features monsemantically. This mixing of correlated features into SAE latents affects both positively and negatively correlated features, meaning that in low L0 SAEs, nearly all latents are both less interpretable and more noisy than an SAE with a correctly set L0.
@@
Finally, we develop a proxy metric based on projections between the SAE decoder and training activations that can detect if L0 is too low. We validate these findings on Gemma-2-2b (gemmateam2024gemma2improvingopen), demonstrating that decoder patterns similar to what we observe in our toy model experiments also manifests in LLM SAEs. We further validate that the optimal L0 we find with our method in Gemma-2-2b matches peak performance on sparse probing tasks (kantamneni2025sparse).
@@ 2 Background
An SAE decomposes an input activation $\mathbf{x}\in\mathbb{R}^{d}$ into a hidden state, $\mathbf{a}$, consisting of $h$ hidden neurons, called “latents”. An SAE is composed of an encoder $\mathbf{W}_{\text{\text{enc}}}\in\mathbb{R}^{h\times d}$, a decoder $\mathbf{W}_{\text{\text{dec}}}\in\mathbb{R}^{d\times h}$, a decoder bias $\mathbf{b}_{\text{\text{dec}}}\in\mathbb{R}^{d}$, and encoder bias $\mathbf{b}_{\text{\text{enc}}}\in\mathbb{R}^{h}$, and a nonlinearity $\sigma$, typically ReLU or a variant like JumpReLU (rajamanoharan2024jumping), TopK (gao2024scaling) or BatchTopK (bussmann2024batchtopk). The decoder is sometimes called the dictionary, in reference to sparse dictionary learning. We use both terms interchangeably.
@@
In this work we focus on BatchTopK and JumpReLU SAEs as these are both considered SOTA architectures. The JumpReLU activation is a modified ReLU with a threshold parameter $\tau>0$, so $\text{JumpReLU}_{\tau}(x)=x\cdot\mathbf{1}_{x>\tau}$. The BatchTopK activation function selects the top $b\times k$ activations across a batch of size $b$, allowing variance in the $k$ selected per sample in the batch. After training, a BatchTopK SAE is converted to a JumpReLU SAE with a global $\tau$. We follow the JumpReLU training procedure outlined by Anthropic (conerly2025jumprelu).
@@ 3 Toy model experiments
The Linear Representation Hypothesis (LRH) (elhage2022toy; park2024lrh) states that LLMs represent concepts (alternatively referred to as “features”) as (nearly) orthogonal linear directions in representation space. Thus, the hidden activations in an LLM are simply the sum all the firing feature vectors (a feature direction with a positive, non-zero magnitude) that are being represented. While an LLM can represent a potentially large number of concepts this way, in any given activation, only a small number of concepts are actively represented.
@@
In these toy model experiments, we mainly focus on BatchTopK SAEs (bussmann2024batchtopk) as this enables direct control of L0. Additionally, we validate our results with JumpReLU SAEs. We train SAEs on 15M synthetic samples with batch size 500 using SAELens (bloom2024saetrainingcodebase).
@@ 3.4 The sparsity–reconstruction tradeoff
It is common practice to evaluate SAE architectures using a sparsity–reconstruction tradeoff plot (huben2024sparse; gao2024scaling; rajamanoharan2024jumping), where the assumption is that having better reconstruction at a given sparsity is inherently better, and indicates that the SAE is correct. Afterall, we train SAEs to reconstruct inputs, so surely an SAE that has better reconstruction must therefore be a better SAE than one that has lower reconstruction?
Sadly, this is not the case. As we discussed in Section 3.3, when the L0 of the SAE is lower than optimal, the SAE can find ways to “cheat” by engaging in feature hedging (chanin2025feature), and get a better MSE score by mixing components of correlated features together. This results in an SAE where the latents are not monosemantic, and do not track ground-truth features.
@@
Figure 5: SAE decoder cosine similarity with true features for the SAEs from Figure 4 with L0=1 (left) and L0=5 (middle), compared with the ground-truth SAE (right). The trained SAEs score much better than the ground truth SAE on variance explained, despite their corrupted, polysemantic latents.
@@ 3.5 Detecting the true L0 using the SAE decoder
Figure 1 reveals that the SAE decoder latents contain mixes of underlying features, both when the L0 is too high and also when it is too low. As the SAE approaches the correct L0, each SAE latent has fewer components of multiple true features mixed in, becoming more monosemantic. Thus, we expect that the closer the SAE is to the correct L0, the more latents should be orthogonal relative to each other, as there are fewer components of shared correlated features mixed into latents. If we are far from the correct L0, then SAE latents contain components of many underlying features, and thus we expect latents to have higher cosine similarity with each other.
We call this metric decoder pairwise cosine similarity, $c_{\text{dec}}$, and define it as below:
$c_{\text{dec}}=\frac{1}{\binom{h}{2}}\sum_{i=1}^{h-1}\sum_{j=i+1}^{h}|\cos(\mathbf{W}_{\text{dec},i},\mathbf{W}_{\text{dec},j})|$ (4)
where $\binom{h}{2}=\frac{h(h-1)}{2}$ is the total number of distinct pairs of latents in the SAE decoder.
If SAE decoder latents are mixing lots of positive and negative components of correlated and anti-correlated features, then each SAE latent should become less orthogonal to each other SAE latent, as many latents will likely mix together similar features. This should mean that the absolute value of the cosine similarity between arbitrary latents should also increase the worse this mixing becomes.
We calculate pairwise calculate similarity $c_{\text{dec}}$ for each of the BatchTopK SAEs we trained on toy models from Section 3.5. Results are shown in Figure 6. We see that pairwise cosine similarity is minimized at the true L0.
Figure 6: Decoder pairwise cosine similarity $c_{\text{dec}}$ evaluated on 5 seeds of toy model SAEs. The true L0 is indicated with a dotted line at 11. Shaded area is 1 stdev. $c_{\text{dec}}$ is minimized at the true L0.
We explore alternative metrics in Appendix A.9. Further toy model experiments are shown in Appendix A.4. Pytorch code implementing $c_{\text{dec}}$ is provided in Appendix A.17. We provide formal theoretical justification for the $c_{\text{dec}}$ metric in Appendix A.6.
@@ 3.6 JumpReLU SAE experiments
So far, we have only investigated BatchTopK SAES due to their ease of setting L0. We now validate that these same conclusions apply to JumpReLU SAEs. We train JumpReLU saes with a range of $\lambda_{s}$ to control the sparsity of the SAEs. We show plots of $\lambda_{s}$ vs L0 and decoder pairwise cosine similarity vs L0 for these SAEs in Figure 7. We see that the cosine similarity vs L0 broadly follows the same pattern as we saw for BatchTopK SAEs, and is minimized at the correct L0.
Figure 7: (left) L0 coefficient $\lambda_{s}$ vs L0 for JumpReLU SAEs. (right) Decoder pairwise cosine similarity vs L0 for JumpReLU SAEs. The true L0, 11, is marked by a dotted line on the plot.
Interestingly, we see that the L0 does not change linearly with $\lambda_{s}$, but instead “sticks” near the correct L0. This is a testament to Anthropic’s JumpReLU SAE training method (conerly2025jumprelu), as a wide range of sparsity coefficients $\lambda_{s}$ cause the SAE to naturally find the correct L0.
@@ 4 LLM experiments
We train a series of BatchTopK SAEs (bussmann2024batchtopk) with $h=32768$ on Gemma-2-2b (gemmateam2024gemma2improvingopen) and Llama-3.2-1b (dubey2024llama3herdmodels) varying L0 and calculate $c_{\text{dec}}$. Each SAE is trained on 500M tokens from the Pile (pile) using SAELens (bloom2024saetrainingcodebase). We also calculate k-sparse probing performance for these SAEs using the benchmark from kantamneni2025sparse, consisting of over 100 sparse probing tasks. Results are shown in Figure 8.
Figure 8: Decoder pairwise cosine similarity vs SAE L0 and K-sparse probing F1 vs L0 with 3 seeds per L0. (left) Gemma-2-2b layer 5 BatchTopK results. (right) Llama-3.2-1b layer 7 BatchTopK SAEs. In both cases, peak sparse probing performance occurs in the elbow just before $c_{\text{dec}}$ jumps due to low L0, although the shapes of the $c_{\text{dec}}$ plots vary at high L0.
The Llama SAE $c_{\text{dec}}$ plot looks very similar to the toy model, with a clear minimum point. The Gemma-2-2b layer 5 SAEs also show a sharp increase in $c_{\text{dec}}$ at low L0 as we saw in toy models, but has a long shallow region with the global minimum actually appearing in that shallow region. In both cases, the “elbow” in the $c_{\text{dec}}$ plots just before the jump due to low L0 is around L0 200, and this also corresponds to peak sparse probing performance. More plots and analysis of $c_{\text{dec}}$ curves are shown in Appendix A.15.
Figure 9: Gemma-2-2b layer 12, with (left) decoder pairwise cosine similarity and K-sparse probing F1 for BatchTopK and JumpReLU SAEs, and (right) normalized decoder projection histograms for BatchTopK SAEs. Histograms are truncated to -20 and 40 to highlight projections near the origin.
@@ 4.1 JumpReLU vs BatchTopK SAEs
We next explore how JumpReLU and BatchTopK SAEs compare with decoder pairwise cosine similarity plots. We train a suite of SAEs on 1B tokens on Gemma-2-2b layer 12. We plot $c_{\text{dec}}$ for a range N values as well as k-sparse probing results for JumpReLU and BatchTopK SAEs in Figure 9 (left).
JumpReLU and BatchTopK SAEs behave similarly at low L0, with the high $c_{\text{dec}}$ at low L0 corresponding to poor sparse-probing performance. However, we see notable differences at high L0. The BatchTopK SAEs have a global $c_{\text{dec}}$ minimum around 200, but JumpReLU SAEs $c_{\text{dec}}$ minimum appears closer to 250-300. As we saw in Figure 8 as well, using the “elbow” of the plots just before $c_{\text{dec}}$ jumps due to low L0 seems to roughly correspond to peak k-sparse probing performance.
For JumpReLU SAEs, we see that $c_{\text{dec}}$ rises much less than BatchTopK SAEs at high L0, and indeed, JumpReLU SAEs also perform much better than BatchTopK SAEs at sparse probing when L0 is high. We suspect this is due to JumpReLU SAEs being able to “stick” near the correct threshold per latent like we saw in our toy models section. We investigate the differences in learned SAEs between JumpReLU and BatchTopK further in Appendix A.16.